OpenStudy (anonymous):

Evaluate the following limit: lim (1+7/x)^x/12 x→∞

OpenStudy (anonymous):

I got the number 0.583

OpenStudy (anonymous):

the correct answer is 1.79. I'm not sure where i went wrong

hartnn (hartnn):

how did you get 0.583 ? mind showing your work/steps ?

OpenStudy (mathmate):

is it \( \large \frac{(1+7/x)^x}{12}\) or \( \large (1+7/x)^{x/12} \)

OpenStudy (anonymous):

y=lim (1+7/x)^x/12 ln(y)=lim x/12 ln(1+7/x) \[\frac{ \ln(1+\frac{ 7 }{x} }{\frac{12 }x }\]

OpenStudy (anonymous):

mathmate it's the second

OpenStudy (abb0t):

\[\lim_{x \rightarrow ∞} \frac{ (1+7)^x }{ 12 }\]

OpenStudy (anonymous):

then i took d/dx to both the numerator and the denominator

OpenStudy (anonymous):

the problem is \[\lim_{x \rightarrow \infty} (1+\frac{ 7 }{ 12})^\frac{ x }{12 }\]

OpenStudy (abb0t):

In order to use L'hopitals rule, you must have a fraction with a function on the numerator and denominator

hartnn (hartnn):

do you know a general formula \[\lim_{y \rightarrow 0} (1+y)^{\frac{ 1}{y}}=...?\]

hartnn (hartnn):

if you know ^ that, then you can put 7/x = y in your limit question first, to bring in that form.

hartnn (hartnn):

if you use the formula, you get the correct answer in just few steps by adjusting the exponent. \[\lim_{y \rightarrow 0} (1+y)^{\frac{ 1}{y}}=e\]

OpenStudy (anonymous):

im trying that right now

hartnn (hartnn):

okay :) take your time...

OpenStudy (anonymous):

thanks i got the right answer

hartnn (hartnn):

good! you're welcome ^_^